OpenStudy (anonymous):

It takes Shonda 3/4 of an hour to dry one load of laundry in her dryer. How many hours will it take her to dry 7 loads of laundry? a 5 1/4 b 3/28 c 2 1/2 d 1 3/4

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

@gabbyalicorn

OpenStudy (gabbyalicorn):

Sorry, im not really good at math :/

OpenStudy (anonymous):

ok

OpenStudy (acxbox22):

multiply 3/4 by 7

OpenStudy (dancing_stars23):

5 1/4 because 3/4 = .75 .75*7= 5.25 which equals 21/4 = 5 1/4 honestly if you got to mathway, they help a lot

OpenStudy (anonymous):

What kind of a name is Shonda? Math textbooks always have the weirdest names... It's like a combination of Honda and Shawn or something.

OpenStudy (dancing_stars23):

@sfaigan lmao

OpenStudy (anonymous):

Honda's newest vehicle. The Shonda 2015. It dries your laundry in 45 minutes.